OpenStudy (anonymous):

One of my answers for a parabola (I'm writing the equation in standard form) came out as a decimal.. Do I write it as a decimal, a fraction, or does it matter? Because I know some answers won't be accepted in decimal form and I can't remember if that applies to Quadratic equations..

OpenStudy (amistre64):

that seems more like a formatting issue with your program and not really a set rule in mathematics

OpenStudy (amistre64):

2.5 = 3/2 = 2 1/2 = .... are all mathically the same

OpenStudy (amistre64):

lol, make that 5/2

OpenStudy (anonymous):

okay lol so it doesn't really matter which I go with because they all come out the same?

OpenStudy (amistre64):

correct, but if the program you are using specifies a format ... then it will only accept that one. When I have to use MathLab i mostly get wrong results for not formatting it to their specifications :/

OpenStudy (amistre64):

what was the problem that created the decimal?

OpenStudy (anonymous):

I created a parabola and was writing out the equation, and I got confused on which way I was supposed to write it lol